OpenStudy (anonymous):

At the canteen, stickers sell for 3 cents each or 4 for 10 cents. Cindy bought some stickers and paid 39 cents. How many stickers did she buy? How would I solve a problem like this? Thank you!

OpenStudy (lgbasallote):

if you divide 39 by 10 you get 3 with remainder 9 now if you divide 9 by 3 then you get 3 so that means...she bought 3 of the 4 for 10 cents and 3 for the one for 3 cents that means she had (4 x 3) + (1 x 3) = 12 + 3 = 15 stickers does that make sense?

OpenStudy (anonymous):

Yes, thank you very much!

OpenStudy (lgbasallote):

welcome